OpenStudy (anonymous):

Help please! The volume of a rectangular prism is 105 cubic centimeters. The length is 7 centimeters and the height is 5 centimeters. What is the width? A.21 cm B.9 cm C.15 cm D.3 cm

OpenStudy (anonymous):

do you know the formula for a rectangular prism?

OpenStudy (anonymous):

V=B*H*H i think

OpenStudy (anonymous):

** B*H*W

OpenStudy (anonymous):

H x L x W....

OpenStudy (anonymous):

oh lol ok V=H*L*W thanks :)

OpenStudy (anonymous):

ya so insert numbers, and make sure the number/variable you are looking for is by itself then solve

OpenStudy (anonymous):

i have the V but dont have the W

OpenStudy (anonymous):

yes so the equation you need to use is \[\frac{ v }{ l \times w}\]

OpenStudy (anonymous):

insert numbers and solve then you should get your answer